Swifty Corporation purchased a new machine for its assembly process on August 1, 2020. The cost of this machine was $125,100. The company estimated that the machine would have a salvage value of $17,100 at the end of its service life. Its life is estimated at 5 years, and its working hours are estimated at 20,000 hours. Year-end is December 31. Compute the depreciation expense under the following methods. (c) Sum-of-the-years'-digits for 2021 (d) Double-declining-balance for 2021
Answer #1

compute depreciation expense

c) Sum of the year digit =108000*5/15*7/12+108000*4/15*5/12 = $33000

d) Double declining balance = 125100*40%*7/12+125100*60%*40%*5/12 = 41700